# Stochastic SIR model with individual heterogeneity and infection-age dependent infectivity on large non-homogeneous random graphs

### v1

@@ 5.1. Convergence of $(\bar{\mu}^{S,N}_{\cdot},\overline{\mathfrak{F}}^{N}(\cdot))$
$\langle\widetilde{\mu}^{S,N}_{t},\varphi\rangle=\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\mathbbm{1}_{\left\{\widetilde{D}^{N}_{i}(t^{-})=0\right\}}\varphi(X_{i}^{N})\,.$ (5.2)
@@
¯$\overline{}\mathfrak{D}^{N}(t):=\mathbb{E}_{0}^{N}\Bigg{[}\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\sup_{r\in[0,t]}|D^{N}_{i}(r)-\widetilde{D}^{N}_{i}(r)|\Bigg{]}\,,$ (5.6)
@@
¯$\overline{}\mathfrak{A}_{i}^{N}(t)=\sum_{j\in\mathcal{S}^{N}(0)}\omega^{N}(i,j)\cdot\big{[}\lambda^{N}_{j}(t-\tau_{j}^{N})-\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})\big{]},$ (5.7)
where $\widetilde{\tau}_{j}^{N}$ is the first jump time corresponding to the process $(\widetilde{D}^{N}_{j}(s))_{s>0}$ in (5.1), so that ¯$\overline{}\mathfrak{A}_{i}^{N}(t)$ captures the discrepancies between the jumps of $(D^{N}_{j})$ and those of $(\widetilde{D}^{N}_{j})$.
@@
that is the subset of individuals infected by time $t$, while being affected by the mean-field infection rate. For any $j\in\widetilde{\mathcal{D}}^{N}(t)$, $\widetilde{A}_{j}^{N}(t)=t-\widetilde{\tau}_{j}^{N}$ can be interpreted as the corresponding infection age, of individual $j$ at time $t$, like $A_{j}^{N}(t)=t-\tau_{j}^{N}$ for any $j\in\mathcal{D}^{N}(t)$. Possibly $\widetilde{A}_{j}^{N}(t)>\eta_{j}^{N}$, thus the individual $j$ has recovered by time $t$ and $\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})=0$. For any $j\in\mathcal{S}^{N}_{0}\setminus\widetilde{\mathcal{D}}^{N}(t)$ on the other hand, $t-\widetilde{\tau}_{j}^{N}<0$, which entails also $\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})=0$. Similar observations hold for the value of $\lambda^{N}_{j}(t-\tau_{j}^{N})$ depending on whether $j\in\mathcal{D}^{N}(t)$ or not, and if yes whether $A_{j}^{N}(t)>\eta_{j}^{N}$ holds or not. It justifies the statement that ¯$\overline{}\mathfrak{A}_{i}^{N}(t)$ corresponds to the component due to the discrepancies between infection ages.
@@
¯$\overline{}\mathfrak{V}_{i}^{N,1}(t)=\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\Big{[}N\omega^{N}(i,j)\cdot\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})-\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})\cdot\mathbb{E}_{0}^{N}[\bar{\lambda}(X_{j}^{N},t-\widetilde{\tau}_{j}^{N})]\Big{]},$ (5.9)
so that ¯$\overline{}\mathfrak{V}_{i}^{N,1}$ concerns the approximation of the transmission rate by its average, where the expectation in the last term averages the randomness of the infection time $\widetilde{\tau}_{j}^{N}$ with the exterior field $\overline{\mathfrak{F}}(.,X_{j}^{N})$ acting on individual $j$.
@@
¯$\overline{}\mathfrak{L}^{N,1}(t,x)=\int_{\mathbf{\mathbb{X}}}\Big{[}\bar{\omega}^{N}(x,x^{\prime})-\bar{\omega}(x,x^{\prime})\Big{]}\cdot\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\bar{\mu}^{S,N}_{0}(\mathrm{d}x^{\prime})\;,$ (5.10)
@@
Thus, ¯$\overline{}\mathfrak{L}^{N,1}$ concerns the approximation of $\bar{\omega}^{N}$ by the kernel $\bar{\omega}$.
@@
¯$\overline{}\mathfrak{E}^{N,1}(t,x)=\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\cdot\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\;\Big{[}\bar{\mu}^{S,N}_{0}-\bar{\mu}^{S}_{0}\Big{]}(\mathrm{d}x^{\prime}).$ (5.12)
We will see in Remark 5.5 that ¯$\overline{}\mathfrak{E}^{N}(t,X^{N}_{i})$ can also be related to the approximation of the empirical measure process $(\widetilde{\mu}^{S,N}_{s})$ by the limiting $(\bar{\mu}^{S}_{s})$.
@@
¯$\overline{}\mathfrak{V}_{i}^{N,0}(t)=\frac{1}{N}\sum_{j\in\mathcal{I}^{N}(0)}\Big{[}N\omega^{N}(i,j)\lambda^{N}_{j}(A_{j}^{N}(0)+t)-\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})\frac{\bar{\lambda}(X_{j}^{N},A_{j}^{N}(0)+t)}{F^{c}_{X_{j}^{N}}(A_{j}^{N}(0))}\Big{]}\,,$ (5.13)
in a similar way as ¯$\overline{}\mathfrak{V}_{i}^{N}(t)$ for the initially infected individuals. We recall from Assumption 2.6 that for any $j\in\mathcal{I}^{N}(0)$ the value of the conditional expectation $\bar{\lambda}(X_{j}^{N},A_{j}^{N}(0)+t)$ is amplified by the denominator $F^{c}_{X_{j}^{N}}(A_{j}^{N}(0))$ to account for the bias in $\lambda^{N}_{j}(A_{j}^{N}(0)+t)$ due to the fact that individual $j$ has not recovered by time 00. The next terms ¯$\overline{}\mathfrak{L}^{N,0}$ and ¯$\overline{}\mathfrak{E}^{N,0}$ are similarly the analogs of respectively ¯$\overline{}\mathfrak{L}^{N,1}$ and ¯$\overline{}\mathfrak{E}^{N,1}$ for the initially infected individuals. For any $x\in\mathbf{\mathbb{X}}$ and $t\geq 0$,
¯$\overline{}\mathfrak{L}^{N,0}(t,x)=\int_{\mathbf{\mathbb{X}}}\int_{0}^{\infty}\Big{[}\bar{\omega}^{N}(x,x^{\prime})-\bar{\omega}(x,x^{\prime})]\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\;,$ (5.14)
@@
¯$\overline{}\mathfrak{E}^{N,0}(t,x)=\int_{\mathbf{\mathbb{X}}}\int_{0}^{\infty}\bar{\omega}(x,x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\Big{[}\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})-\bar{\mu}^{I}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\Big{]}\,,$ (5.15)
so that ¯$\overline{}\mathfrak{E}^{N,0}(t,X^{N}_{i})$ accounts for the approximation of the initial condition $\bar{\mu}^{I,N}_{0}$ by $\bar{\mu}^{I}_{0}$.
@@
$\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}(s,x^{\prime})\bar{\mu}^{S}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s=\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})\,.$ (5.16)
@@
$\overline{\mathfrak{F}}^{0}(t,x)=\int_{\mathbf{\mathbb{X}}}\int_{0}^{\infty}\bar{\omega}(x,x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\,,$ (5.17)
@@
$\begin{split}\overline{\mathfrak{F}}_{i}^{N,0}(t)-\overline{}\mathfrak{V}_{i}^{N,0}(t)&=\frac{1}{N}\sum_{j\in\mathcal{I}^{N}(0)}\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})\frac{\bar{\lambda}(X_{j}^{N},A_{j}^{N}(0)+t)}{F^{c}_{X_{j}^{N}}(A_{j}^{N}(0))}\\ &=\int_{\mathbf{\mathbb{X}}}\int_{0}^{\infty}\bar{\omega}^{N}(X_{i}^{N},x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\,.\end{split}$
@@
$\begin{split}\overline{\mathfrak{F}}_{i}^{N,0}(t)-\overline{}\mathfrak{V}_{i}^{N,0}(t)-\overline{}\mathfrak{L}^{N,0}(t,X^{N}_{i})&=\int_{\mathbf{\mathbb{X}}}\int_{0}^{\infty}\bar{\omega}(X^{N}_{i},x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\\ &=\overline{}\mathfrak{E}^{N,0}(t,X^{N}_{i})+\overline{\mathfrak{F}}^{0}(t,X_{i}^{N})\,,\end{split}$ (5.21)
which concludes our first claim. For the second claim, there is a first additional step where $A_{i}^{N}(t)$ is related to $\widetilde{A}_{i}^{N}(t)=t-\widetilde{\tau}_{i}^{N}$, through ¯$\overline{}\mathfrak{A}_{i}^{N}(t)$, before we can exploit the same arguments. We first combine (5.7) with (5.20), and hence,
@@
In the expression of ¯$\overline{}\mathfrak{E}^{N,1}(t,x)$ in (5.12), we decided to relate as directly as possible to the difference between $\bar{\mu}^{S,N}_{0}$ and $\bar{\mu}^{S}_{0}$. That being said, this term has the following alternative interpretation in terms of the difference between the processes $\widetilde{\mu}^{S,N}_{s}$ and $\bar{\mu}^{S}_{s}$:
¯$\overline{}\mathfrak{E}^{N,1}(t,x)=\mathbb{E}_{0}^{N}\Bigg{[}\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\cdot\overline{\mathfrak{F}}(s,x^{\prime})\Big{[}\widetilde{\mu}^{S,N}_{s}-\bar{\mu}^{S}_{s}\Big{]}(\mathrm{d}x^{\prime})\;\mathrm{d}s\,\Bigg{]}\,.$ (5.24)
@@
In the expression in the second line, we can replace $Q_{j}$ by its intensity since the integrant is predictable with respect to its filtration (~$(\widetilde{\mathcal{F}}^{N}_{t})$, then express the sum over $j\in\mathcal{S}^{N}(0)$ in terms of the empirical measure $\widetilde{\mu}^{S,N}_{s}$, recall (5.2), which leads to
@@
In the following, we treat separately the various terms distinguished in Lemma 5.4, first ¯$\overline{}\mathfrak{A}_{i}^{N}$, see Lemma 5.6, second ¯$\overline{}\mathfrak{V}^{N,1}$ and ¯$\overline{}\mathfrak{V}^{N,0}$, see Lemma 5.7, third ¯$\overline{}\mathfrak{L}^{N,1}$ and ¯$\overline{}\mathfrak{L}^{N,0}$, see Lemma 5.8, and finally ¯$\overline{}\mathfrak{E}^{N,1}$ and ¯$\overline{}\mathfrak{E}^{N,0}$, see Lemma 5.10. We start with the processes (¯$(\overline{}\mathfrak{A}_{i}^{N})$ defined in (5.7).
@@
Under Assumptions 2.2, 3.1, 3.3 and 3.5, there exist a constant $C>0$ such that
@@
holds a.s. for any $t\geq 0$, where ¯$\overline{}\mathfrak{D}^{N}(t)$ is defined in (5.6), and ¯$\overline{}\mathfrak{V}^{N}$ is given by
¯$\overline{}\mathfrak{V}^{N}:=\lambda^{*}\cdot\sqrt{\Upsilon^{N}+\omega^{*}\cdot\bar{\gamma}^{N}}\,.$ (5.27)
We recall the defining property of $\omega^{*}$ given just after Lemma 3.4.
@@
¯$\overline{}\mathfrak{U}^{N}:=\frac{\lambda^{*}}{N}\mathbb{E}_{0}^{N}\Bigg{[}\sum_{j\in\mathcal{S}^{N}(0)}\big{|}\widetilde{S}_{j}^{N}\big{|}\Bigg{]}\,.$ (5.31)
So Lemma 5.6 will be concluded by showing that ¯$\overline{}\mathfrak{U}^{N}\leq\overline{}\mathfrak{V}^{N}$. Thanks twice to the Cauchy-Schwartz inequality, with the fact that the cardinality of $\mathcal{S}^{N}(0)$ is less than $N$:
$\begin{split}\mathbb{E}_{0}^{N}\Bigg{[}\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\Big{|}\widetilde{S}_{j}^{N}\Big{|}\Bigg{]}\,&\leq\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\sqrt{\mathbb{E}_{0}^{N}\Bigg{[}\Big{(}\widetilde{S}_{j}^{N}\Big{)}^{2}\Bigg{]}}\\ &\leq\sqrt{\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\mathbb{E}_{0}^{N}\Bigg{[}\Big{(}\widetilde{S}_{j}^{N}\Big{)}^{2}\Bigg{]}}\,.\end{split}$ (5.32)

### v2

@@ 5.1. Convergence of $(\bar{\mu}^{S,N}_{\cdot},\overline{\mathfrak{F}}^{N}(\cdot))$
$\langle\widetilde{\mu}^{S,N}_{t},\varphi\rangle=\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\mathbbm{1}_{\left\{\widetilde{D}^{N}_{i}(t)=0\right\}}\varphi(X_{i}^{N})\,.$ (5.2)
@@
$\overline{\mathfrak{D}}^{N}(t):=\mathbb{E}_{0}^{N}\Bigg[\frac{1}{N}\sum_{i\in\mathcal{S}^{N}(0)}\sup_{r\in[0,t]}|D^{N}_{i}(r)-\widetilde{D}^{N}_{i}(r)|\Bigg]\,$, (5.6)
@@
$\overline{\mathfrak{A}}_{i}^{N}(t)=\sum_{j\in\mathcal{S}^{N}(0)}\omega^{N}(i,j)\cdot\big[\lambda^{N}_{j}(t-\tau_{j}^{N})-\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})\big],$ (5.7)
where $\widetilde{\tau}_{j}^{N}$ is the jump time of the process $(\widetilde{D}^{N}_{j}(s))_{s>0}$ in (5.1), so that $\overline{\mathfrak{A}}_{i}^{N}(t)$ captures the discrepancies between the jump of $(D^{N}_{j})$ and that of $(\widetilde{D}^{N}_{j})$.
@@
that is the subset of individuals infected by time $t$, while being affected by the mean-field infection rate. For any $j\in\widetilde{\mathcal{D}}^{N}(t)$, $\widetilde{A}_{j}^{N}(t)=t-\widetilde{\tau}_{j}^{N}$ can be interpreted as the corresponding infection age, of individual $j$ at time $t$, like $A_{j}^{N}(t)=t-\tau_{j}^{N}$ for any $j\in\mathcal{D}^{N}(t)$. Possibly $\widetilde{A}_{j}^{N}(t)>\eta_{j}^{N}$, thus the individual $j$ has recovered by time $t$ and $\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})=0$. For any $j\in\mathcal{S}^{N}_{0}\setminus\widetilde{\mathcal{D}}^{N}(t)$ on the other hand, $t-\widetilde{\tau}_{j}^{N}<0$, which entails also $\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})=0$. Similar observations hold for the value of $\lambda^{N}_{j}(t-\tau_{j}^{N})$ depending on whether $j\in\mathcal{D}^{N}(t)$ or not, and if yes whether $A_{j}^{N}(t)>\eta_{j}^{N}$ holds or not. It justifies the statement that $\overline{\mathfrak{A}}_{i}^{N}(t)$ corresponds to the component due to the discrepancies between infection ages.
@@
$\overline{\mathfrak{V}}_{i}^{N,1}(t)=\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\Big[N\omega^{N}(i,j)\cdot\lambda^{N}_{j}(t-\widetilde{\tau}_{j}^{N})-\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})\cdot\mathbb{E}_{0}^{N}[\bar{\lambda}(X_{j}^{N},t-\widetilde{\tau}_{j}^{N})]\Big],$ (5.9)
so that $\overline{\mathfrak{V}}_{i}^{N,1}$ concerns the approximation of the transmission rate by its average, where the expectation in the last term averages the randomness of the infection time $\widetilde{\tau}_{j}^{N}$ with the exterior field $\overline{\mathfrak{F}}(.,X_{j}^{N})$ acting on individual $j$.
@@
$\overline{\mathfrak{L}}^{N,1}(t,x)=\int_{\mathbf{\mathbb{X}}}\Big[\bar{\omega}^{N}(x,x^{\prime})-\bar{\omega}(x,x^{\prime})\Big]\cdot\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\bar{\mu}^{S,N}_{0}(\mathrm{d}x^{\prime})\;,$ (5.10)
@@
Thus, $\overline{\mathfrak{L}}^{N,1}$ concerns the approximation of $\bar{\omega}^{N}$ by the kernel $\bar{\omega}$.
@@
$\overline{\mathfrak{E}}^{N,1}(t,x)=\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\cdot\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\;\Big[\bar{\mu}^{S,N}_{0}-\bar{\mu}^{S}_{0}\Big](\mathrm{d}x^{\prime}).$ (5.12)
We will see in Remark 5.5 that $\overline{\mathfrak{E}}^{N}(t,X^{N}_{i})$ can also be related to the approximation of the empirical measure process $(\widetilde{\mu}^{S,N}_{s})$ by the limiting $(\bar{\mu}^{S}_{s})$.
@@
$\overline{\mathfrak{V}}_{i}^{N,0}(t)=\frac{1}{N}\sum_{j\in\mathcal{I}^{N}(0)}\Big[N\omega^{N}(i,j)\lambda^{N}_{j}(A_{j}^{N}(0)+t)-\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})\frac{\bar{\lambda}(X_{j}^{N},A_{j}^{N}(0)+t)}{F^{c}_{X_{j}^{N}}(A_{j}^{N}(0))}\Big]\,$, (5.13)
in a similar way as $\overline{\mathfrak{V}}_{i}^{N,1}(t)$, now for the initially infected individuals. We recall from Assumption 2.6 that for any $j\in\mathcal{I}^{N}(0)$ the value of the conditional expectation $\bar{\lambda}(X_{j}^{N},A_{j}^{N}(0)+t)$ is amplified by the denominator $F^{c}_{X_{j}^{N}}(A_{j}^{N}(0))$ to account for the bias in $\lambda^{N}_{j}(A_{j}^{N}(0)+t)$ due to the fact that individual $j$ has not recovered by time 0. The next terms $\overline{\mathfrak{L}}^{N,0}$ and $\overline{\mathfrak{E}}^{N,0}$ are similarly the analogs of respectively $\overline{\mathfrak{L}}^{N,1}$ and $\overline{\mathfrak{E}}^{N,1}$, now for the initially infected individuals. For any $x\in\mathbf{\mathbb{X}}$ and $t\geq 0$,
$\overline{\mathfrak{L}}^{N,0}(t,x)=\int_{\mathbf{\mathbb{X}}}\hskip-2.84544pt\int_{0}^{\infty}\Big[\bar{\omega}^{N}(x,x^{\prime})-\bar{\omega}(x,x^{\prime})]\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\;,$ (5.14)
@@
$\overline{\mathfrak{E}}^{N,0}(t,x)=\int_{\mathbf{\mathbb{X}}}\hskip-2.84544pt\int_{0}^{\infty}\bar{\omega}(x,x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\Big[\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})-\bar{\mu}^{I}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\Big]\,$, (5.15)
so that $\overline{\mathfrak{E}}^{N,0}(t,X^{N}_{i})$ accounts for the approximation of the initial condition $\bar{\mu}^{I,N}_{0}$ by $\bar{\mu}^{I}_{0}$.
@@
$\int_{0}^{t}\hskip-5.69046pt\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\overline{\mathfrak{F}}(s,x^{\prime})\bar{\mu}^{S}_{s}(\mathrm{d}x^{\prime})\mathrm{d}s=\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\mathbb{E}[\bar{\lambda}(x^{\prime},t-\widetilde{\tau}_{x^{\prime}})]\bar{\mu}^{S}_{0}(\mathrm{d}x^{\prime})\,.$ (5.16)
@@
$\overline{\mathfrak{F}}^{0}(t,x)=\int_{\mathbf{\mathbb{X}}}\hskip-2.84544pt\int_{0}^{\infty}\bar{\omega}(x,x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\bar{\mu}_{0}^{I}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\,$, (5.17)
@@
$\begin{split}\overline{\mathfrak{F}}_{i}^{N,0}(t)-\overline{\mathfrak{V}}_{i}^{N,0}(t)&=\frac{1}{N}\sum_{j\in\mathcal{I}^{N}(0)}\bar{\omega}^{N}(X_{i}^{N},X_{j}^{N})\frac{\bar{\lambda}(X_{j}^{N},A_{j}^{N}(0)+t)}{F^{c}_{X_{j}^{N}}(A_{j}^{N}(0))}\\ &=\int_{\mathbf{\mathbb{X}}}\hskip-2.84544pt\int_{0}^{\infty}\bar{\omega}^{N}(X_{i}^{N},x^{\prime})\cdot\frac{\bar{\lambda}(x^{\prime},a^{\prime}+t)}{F^{c}_{x^{\prime}}(a^{\prime})}\;\bar{\mu}^{I,N}_{0}(\mathrm{d}x^{\prime},\mathrm{d}a^{\prime})\,.\end{split}$
@@
[…]
which concludes our first claim. For the second claim, there is a first additional step where $A_{i}^{N}(t)$ is related to $\widetilde{A}_{i}^{N}(t)=t-\widetilde{\tau}_{i}^{N}$, through $\overline{\mathfrak{A}}_{i}^{N}(t)$, before we can exploit the same arguments. We first combine (5.7) with (5.20), and hence,
@@
In the expression of $\overline{\mathfrak{E}}^{N,1}(t,x)$ in (5.12), we decided to relate as directly as possible to the difference between $\bar{\mu}^{S,N}_{0}$ and $\bar{\mu}^{S}_{0}$. That being said, this term has the following alternative interpretation in terms of the difference between the processes $\widetilde{\mu}^{S,N}_{s}$ and $\bar{\mu}^{S}_{s}$:
$\overline{\mathfrak{E}}^{N,1}(t,x)=\mathbb{E}_{0}^{N}\Bigg[\int_{0}^{t}\int_{\mathbf{\mathbb{X}}}\bar{\omega}(x,x^{\prime})\bar{\lambda}(x^{\prime},t-s)\cdot\overline{\mathfrak{F}}(s,x^{\prime})\Big[\widetilde{\mu}^{S,N}_{s}-\bar{\mu}^{S}_{s}\Big](\mathrm{d}x^{\prime})\;\mathrm{d}s\,\Bigg]\,.$ (5.24)
@@
In the expression in the second line, we can replace $Q_{j}$ by its intensity since the integrant is predictable with respect to its filtration $(\widetilde{\mathcal{F}}^{N}_{t})$, then express the sum over $j\in\mathcal{S}^{N}(0)$ in terms of the empirical measure $\widetilde{\mu}^{S,N}_{s}$, recall (5.2), which leads to
@@
In the following, we treat separately the various terms distinguished in Lemma 5.4, first $\overline{\mathfrak{A}}_{i}^{N}$, see Lemma 5.6, second $\overline{\mathfrak{V}}^{N,1}$ and $\overline{\mathfrak{V}}^{N,0}$, see Lemma 5.7, third $\overline{\mathfrak{L}}^{N,1}$ and $\overline{\mathfrak{L}}^{N,0}$, see Lemma 5.8, and finally $\overline{\mathfrak{E}}^{N,1}$ and $\overline{\mathfrak{E}}^{N,0}$, see Lemma 5.10. We start with the processes $(\overline{\mathfrak{A}}_{i}^{N})$ defined in (5.7).
@@
Under Assumptions 2.2, 3.1, 3.2 and 3.4, there exist a constant $C>0$ such that
@@
holds a.s. for any $t\geq 0$, where $\overline{\mathfrak{D}}^{N}(t)$ is defined in (5.6), and $\overline{\mathfrak{V}}^{N}$ is given by
$\overline{\mathfrak{V}}^{N}:=\lambda^{*}\cdot\sqrt{\Upsilon^{N}+\omega^{*}\cdot\bar{\gamma}^{N}}\,.$ (5.27)
We recall the defining property of $\omega^{*}$ given in Assumption 3.3.
@@
$\overline{\mathfrak{U}}^{N}:=\frac{\lambda^{*}}{N}\mathbb{E}_{0}^{N}\Bigg[\sum_{j\in\mathcal{S}^{N}(0)}\big|\widetilde{S}_{j}^{N}\big|\Bigg]\,.$ (5.31)
So Lemma 5.6 will be concluded by showing that $\overline{\mathfrak{U}}^{N}\leq\overline{\mathfrak{V}}^{N}$. Thanks to the Cauchy-Schwartz inequality with respect to the product measure, with the fact that the cardinality of $\mathcal{S}^{N}(0)$ is less than $N$:
$\begin{split}\mathbb{E}_{0}^{N}\Bigg[\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\Big|\widetilde{S}_{j}^{N}\Big|\Bigg]\,\leq\sqrt{\frac{1}{N}\sum_{j\in\mathcal{S}^{N}(0)}\mathbb{E}_{0}^{N}\Bigg[\Big(\widetilde{S}_{j}^{N}\Big)^{2}\Bigg]}\,.\end{split}$ (5.32)